Suppose there are 16 balls in a urn. They are identical except for color. 5 balls are red and 4 are blue and 7 are yellow. You are interested to draw out one ball, note its color, and set it aside. Then you are to draw another ball and note its color.
A) what are the outcomes of the experiment?
B) What is the probability of each outcome?

Answer by Edwin McCravy(20054)

Suppose there are 16 balls in a urn. They are identical
except for color. 5 balls are red and 4 are blue and 7
are yellow. You are interested to draw out one ball, note
its color, and set it aside. Then you are to draw another
ball and note its color.
A) what are the outcomes of the experiment?
{RR,RB,RY,BR,BB,BY,YR,YB,YY}
B) What is the probability of each outcome?
P(RR) = (5/16)(4/15) = 1/12 
P(RB) = (5/16)(4/15) = 1/12 
P(RY) = (5/16)(7/15) = 7/48 
P(BR) = (4/16)(5/15) = 1/12 
P(BB) = (4/16)(3/15) = 1/20
P(BY) = (4/16)(7/15) = 7/60
P(YR) = (7/16)(5/15) = 7/48
P(YB) = (7/16)(4/15) = 7/60
P(YY) = (7/16)(6/15) = 7/40

They all add up to 1.  The LCD of them all is 240.
You can add them to show they add up to 1.

Edwin